Knowing that delta H of water is 40.7 Kj/mol

How many KJ will be required to vaporize 50.0 g of water?

Calculate the percent error if you obtain a value of 39.2 Kj/mol for delta h vap for water for experiment
Answer #1

Hvap = 40.7 kJ/mol

m = 50 g of water

mol = mass/MW = 50/18 = 2.7778 mo of water

then

Qtotal = n*Hvap = 2.777 * 40.7

Qtotal = 113.0239 kJ

B)

Erorr %

given

Hvap = 39.2

%error = (Real - Experimental ) / Real * 100% = (40.7 -39.2 )/40.7 * 100 % = 3.685 % under estimation